OpenStudy (anonymous):

If a 2.5 g sample of gold is hammered into an equilateral triangle that is .28mm thick, how long are the sides?

OpenStudy (ineedhelplz):

Okay, let's first determine the volume of the triangle. Volume of the triangle = Mass of the triangle/Density of gold = 2.5/19.3 = 0.129533678cm³ 1mm= 0.1cm 0.28mm= (0.1)(0.28) = 0.028cm Volume of a solid = (area of its cross section)(thickness) Area of the triangle = Volume/thickness = 0.129533678/0.028 = 4.626202813cm²

OpenStudy (anonymous):

I think I know how to do it. I need to find the height or something right? :/

OpenStudy (anonymous):

Oh ok

OpenStudy (anonymous):

that works too

OpenStudy (anonymous):

wait where did you get the 19.3 from?

OpenStudy (ineedhelplz):

Multiplying them

OpenStudy (ineedhelplz):

Now, in an equilateral triangle, each angle is 60°. The area of a triangle is determined using the formula (1/2)absin60°, where a and b are the lengths of two of its sides. Since a=b in the case of the equilateral triangle, its area= (1/2)L²sin60°, where L is the length of each of its sides. (1/2)L²sin60°= 4.626202813 L²= (2)(4.626202813)/sin60° L²= 10.68375776 L= |√10.68375776| = 3.268601805cm = 3.3cm correct to 2 significant figures

OpenStudy (ineedhelplz):

Does that answer your question? :)

OpenStudy (anonymous):

Somewhat....

OpenStudy (anonymous):

How could I figure out the lenght of the sides though?

OpenStudy (anonymous):

Now I'm just confused.. :/

OpenStudy (ineedhelplz):

The answer is = 3.268601805cm

OpenStudy (anonymous):

Really? For all three sides?

OpenStudy (ineedhelplz):

For 2

OpenStudy (ineedhelplz):

actually

OpenStudy (anonymous):

Oh yeah beacuse it is an equilateral triangle which means all sides are the same

OpenStudy (ineedhelplz):

L²= (2)(4.626202813)/sin60° L²= 10.68375776 L= |√10.68375776|

OpenStudy (ineedhelplz):

Those are the sides

OpenStudy (ineedhelplz):

So does it make sense now?

OpenStudy (ineedhelplz):

and yes it is because its equilateral triangle :)

OpenStudy (anonymous):

Yeah! So the two answers are the 10.6 and 3.26?

OpenStudy (ineedhelplz):

All three sides are 10.6 :D GoodJob

OpenStudy (anonymous):

Thank you!

OpenStudy (ineedhelplz):

You're welcome :)